### v1

# Sharing Key Semantics in Transformer Makes Efficient Image Restoration

Abstract
Image Restoration (IR), a classic low-level vision task, has witnessed significant advancements through deep models that effectively model global information. Notably, the Vision Transformers (ViTs) emergence has further propelled these advancements. When computing, the self-attention mechanism, a cornerstone of ViTs, tends to encompass all global cues, even those from semantically unrelated objects or regions. This inclusivity introduces computational inefficiencies, particularly noticeable with high input resolution, as it requires processing irrelevant information, thereby impeding efficiency. Additionally, for IR, it is commonly noted that small segments of a degraded image, particularly those closely aligned semantically, provide particularly relevant information to aid in the restoration process, as they contribute essential contextual cues crucial for accurate reconstruction. To address these challenges, we propose boosting IR’s performance by sharing the key semantics via Transformer for IR (i.e., SemanIR) in this paper. Specifically, SemanIR initially constructs a sparse yet comprehensive key-semantic dictionary within each transformer stage by establishing essential semantic connections for every degraded patch. Subsequently, this dictionary is shared across all subsequent transformer blocks within the same stage. This strategy optimizes attention calculation within each block by focusing exclusively on semantically related components stored in the key-semantic dictionary. As a result, attention calculation achieves linear computational complexity within each window. Extensive experiments across 6 IR tasks confirm the proposed SemanIR’s state-of-the-art performance, quantitatively and qualitatively showcasing advancements.
1 Introduction
Image restoration (IR) stands as a fundamental task within low-level computer vision, aiming to enhance the quality of images affected by numerous factors, including noise, blur, low resolution, compression artifacts, mosaic patterns, adverse weather conditions, and other forms of distortion. This capability holds broad utility across various domains, facilitating information recovery in medical imaging, surveillance, and satellite imagery. Furthermore, it bolsters downstream vision tasks like object detection, recognition, and tracking [79, 65]. Despite notable progress in recent years, prevalent IR methods encounter challenges in effectively addressing complex distortions or preserving/recovering crucial image details [50]. Achieving high-quality image recovery necessitates meticulous exploration of the rich information present in degraded counterparts.
Figure 1: (a) The CNN filter captures information only within a local region. (b) The standard MLP/Transformer architectures take full input in a long-sequence manner. (c) The window-size multi-head self-attention (MSA) mechanism builds a full connection within each window. (d) Position-fixed sparse connection. (e) The proposed Key-Semantic connection.
In modern IR systems, representative networks for learning rich image information are typically constructed using three fundamental architectural paradigms. i.e., convolutional neural networks (CNNs) [44, 98], Multilayer perceptrons (MLPs) [5, 87], and Vision Transformers (ViTs) [89, 21]. The input image is treated as a regular grid of pixels in the Euclidean space for CNNs (Fig. 1(a)) or a sequence of patches for MLPs and ViTs (Fig. 1(b)). However, degraded inputs usually contain irregular and complex objects. While these choices perform admirably in scenarios with regular or well-organized object boundaries, they have limitations when applied to degraded images with more flexible and complex geometrical contexts.
Additionally, CNNs struggle to model long-range dependencies due to their limited receptive field (Fig. 1(a)). In contrast, MLPs and ViTs can capture long-range relations effectively, although at the cost of losing inductive bias and incurring a heavy computational burden, i.e., quadratic complexity increases with higher input resolution[87, 89, 21, 75]. To address these limitations, recent IR methods have explored strategies for complexity reduction. A common approach is to implement MSA within local image regions [57, 52]. For example, SwinIR [52] and GRL [50] employ full MSA or region-fixed anchored stripe MSA, but they still struggle with irregular object connections. Furthermore, prior research [114] highlights that smooth image content is more prevalent than complex image details, underscoring the need for distinct treatment based on semantic content.
In this paper, we introduce a novel approach, SemanIR, to address the limitations above. Specifically, within each transformer stage, we first construct a key-semantic dictionary, which stores only the top-k semantically related relations for each given degraded patch with the k-nearest neighbors (i.e., KNN) algorithm. Then, the attention operation within each transformer layer occurs only among the top-k patches. This design brings two main advantages, i.e., 1) Each degraded patch benefits from its semantically similar patches, typically containing comparable contextual or textural information, while excluding the side effects from other patches that contain entirely unrelated information. 2) Compared to the conventional window-wise attention, which built a dense connection between all the patches (Fig. 1(c)) that leads to highly computationally demanding, or a sparse but position-fixed manner (Fig. 1(d)) which introduces irrelevant semantics. Our key-semantic connection (Fig. 1(e)) leads to a sparse yet more meaningful attention operation, which allows our method to achieve the same receptive field as previous ViTs-based methods while maintaining lower computational costs. This is not like previous token merging or pruning methods [6, 95, 74] that may merge unrelated information or prune some semantically related information. In addition, to make the proposed method more efficient, instead of creating a key-semantic dictionary for each transformer layer, we create it just once at the beginning of each transformer stage and then share it with all the following transformer layers within the same stage. This not only largely reduced the computation burden but also made our methods different from other token pruning and merging methods [110, 93], which include dynamic patch skipping/selection within each attention layer or involve an additional offset generation network. In addition, such a sharing strategy allows each degraded patch can be continuously optimized by its semantically related patches within each stage.
It is also worth noting that the implementation of the attention layer of our SemanIR is achieved in three interesting manners (i.e., Triton [20] [^1] Open-source GPU programming tool https://openai.com/research/triton., torch-mask, and torch-gather), which are discussed in our ablation studies. The contributions of this work are:
  1. 1.
    For each degraded input patch, we propose to construct a key-semantic dictionary that stores its most semantically relevant $k$ patches in a sparse yet representative manner. This strategy excludes the side effects of a given degraded patch from semantically unrelated parts.
  2. 2.
    Based on the constructed key-semantic dictionary, we propose to share the key semantic information across all the attention layers within each transformer stage, which not only makes each degraded patch well-optimized but also largely reduces the computational complexity compared to conventional attention operations.
  3. 3.
    Extensive experimental results show that the proposed SemanIR achieves state-of-the-art performance on 6 IR tasks, i.e., deblurring, JPEG compression artifact removal (JPEG CAR), denoising, IR in adverse weather conditions (AWC), demosaicking, and classic image super-resolution (SR).
2 Related Work
Image Restoration (IR), as a long-standing ill-posed inverse problem, is designed to reconstruct the high-quality image from the corresponding degraded counterpart with numerous applications [77, 3, 51]. Initially, IR was addressed through model-based solutions, involving the search for solutions to specific formulations. However, learning-based approaches have gained much attention with the significant advancements in deep neural networks. Numerous approaches have been developed, including regression-based [53, 43, 52, 15, 50, 108] and generative model-based pipelines [26, 91, 60, 96]. In this paper, we propose a regression-based method for image restoration.
Non-Local Priors Modeling in IR. Tradition model-based IR methods reconstruct the image by regularizing the results (e.g., Tikhonov regularization [28]) with formulaic prior knowledge of natural image distribution. However, it is challenging for these methods to recover realistic detailed results with hand-designed priors. Besides, some other classic method finds that self-similarity is an effective prior, which leads to an impressive performance [7, 18]. Apart from traditional methods, the non-local prior has also been utilized in modern deep learning networks [55, 90, 50, 112], typically captured by the self-attention mechanism. More recently, the overwhelming success of transformers [89] in the natural language processing domain [37] and the classic vision community [21, 8, 86, 94, 10] has led to the development of numerous ViT-based IR methods. These methods aim to enhance the learning ability for modeling non-local priors [52, 97, 50, 14, 16, 102] and consequently archives better performance. Meanwhile, this raises a question: are all non-local priors essential for IR?
Key-Semantic Non-local Prior Exploration for IR. To answer the question, we found many methods demonstrating the effectiveness of modeling key semantics within ViTs. For example, KiT [45] proposed increasing the non-local connectivity between patches at different positions through KNN matching. This approach aims to better capture the non-local relations between the base patch and other patches in each attention calculation. However, it results in significant additional computational costs due to the KNN matching. DRSformer [13] proposed a top-k selection strategy that chooses the most relevant tokens to model the non-local priors for draining after each self-attention calculation without reducing the computation complexity. Similar conclusions can be also drawn from the graph perspective solutions [29, 78, 66, 31, 54, 34] for various IR tasks, like facial expression restoration [58], image denoising [82], and artifact reduction [66]. [34] construct the graph with transformer-based architecture where each patch is connected to all other patches. All these methods suggest that if the semantically related information can be addressed, the degraded image can be restored with better performance. However, the efficiency issue, which is extremely unignorable, remains untouched within the aforementioned methods. It is particularly crucial for ViTs-based image restoration methods, which often need to address high-resolution degraded input images. In this paper, we propose to share key semantics within each transformer stage, an approach validated to be efficient and effective via experimental and theoretical analysis.
Figure 2: The proposed SemanIR mainly consists of a convolutional feature extractor, the main body of SemanIR for representation learning, and an image reconstructor. The main body in columnar shape shown here is for image SR, while the U-shaped structure (shown in Appx. A.2) is used for other IR tasks. (b) The transformer layer of our SemanIR. The toy example of $k$=3 for (c) the Key-semantic dictionary construction and (d) the attention of each Layer.
3 Methodology
To comprehensively study the effectiveness of the proposed method that is architecture-agnostic for various IR tasks, we adopted two of the most commonly used architectures i.e.the columnar architecture (shown in Fig. 2 (a)) for image SR and the U-shaped architecture (shown in the Appendix, i.e., Appx. A.2) for other IR tasks. In the following, we first show how to construct the key-semantic dictionary in Sec. 3.1. Based on the key-semantic dictionary, then we explain why sharing it works for IR, and we introduce the basic unit, the key-semantic transformer layer in Sec. 3.2. Finally, two interesting discussions (Sec.3.3) are introduced regarding the implementation style of the Key-Graph attention and two top-k settings during the training. The efficiency analysis is provided in Appx. A.3.
3.1 Key-Semantic Dictionary Construction
Consider the input feature $F_{in}\in\mathbb{R}^{H\times W\times C}$, where $H$, $W$, and $C$ denote the height, the width, and the channel. ViTs are good at modeling global dependencies for $F_{in}$. This is achieved by the MSA, the core of ViTs, by connecting all other patches to a certain patch. Specifically, $F_{in}$ is first split into $N$ non-overlapping patches, forming its patch representation $\mathcal{P}=\{p_{i}|p_{i}\in\mathbb{R}^{hw\times c},i=1,2,3,...,N\}$, where $h$, $w$, and $c$ are the height, the width, and the channel of each patch. To achieve such global connectivity $\mathcal{D}$, $\mathcal{P}$ is linearly projected into Query ($Q$), Key ($K$), and Value ($V$) matrices, which are denoted as $Q=\mathcal{P}\mathbf{W}_{qry}$, $K=\mathcal{P}\mathbf{W}_{key}$, and $V=\mathcal{P}\mathbf{W}_{val}$. $\mathbf{W}_{qry/key/val}$ represents the learnable projection weights. Then $\mathcal{D}$ is performed by a softmax function as follows:
$\mathcal{D}_{ij}=\frac{\operatorname{exp}(Q_{i}K_{j}^{\top})}{\sum_{k=1...j}\operatorname{exp}(Q_{i}K_{k}^{\top}/\sqrt{d})},i=1,2,3,...,N,$ (1)
where $d$ is the dimension of $Q$ and $K$. Then each patch is aggregated via $\sum_{i}\mathcal{D}_{ij}V_{i}$. However, $\mathcal{D}_{ij}$ functions as a full semantic dictionary, where each patch is connected to all other patches regardless of their semantic relatedness. e.g., given a sub-graph with a green dog patch shown in Fig. 1(c), the tree-related patches are also considered. Since such an operation occurs at each attention calculation step in ViTs, it inevitably increases the computational cost, especially for large-scale inputs. In addition, for IR, a degraded patch usually benefits from its most semantically related patches, as they share similar texture and geometric information. This naturally raises the question: Can we build a key-semantic dictionary, $\mathcal{D}_{K}$, where each patch is connected only to its most related patches?
To mitigate this problem, given $\mathcal{P}$, we first construct a fully connected dictionary $\mathcal{D}$ by calculating its self-similarity $\operatorname{Sim()}$ via a naive dot product operation as $\mathcal{D}(i,j)=\operatorname{Sim}(i,j)=p_{i}\cdot p_{j}^{\top}$, which describes the correlation among all the patches, with higher values indicating stronger correlations. To reduce the side influence of patches with low correlation (e.g., the tree-related patches at the upper left part in Fig. 1 (c)) for the green background dog destination patch, we keep only $k$ highly related patches and exclude the remaining. This is achieved by a KNN algorithm from $\mathcal{D}$ as follows:
$\mathcal{D}_{K}(i,j)=\begin{cases}\mathcal{D}(i,j),~{}\mathcal{D}(i,j)\geq\operatorname{Sim}(i,)_{k}~{}\text{and}~{}i\neq j\\ 0,~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{otherwise},\end{cases}$ (2)
where $\operatorname{Sim}(i,)_{k}$ denotes the $k_{th}$ largest connectivity value of patch $p_{i}$. As a result, $\mathcal{D}_{K}$ contains only the patches with high correlation (e.g., dog-related patches in Fig. 1(e)) for the destination patch (e.g., the green dog patch). We formalize the key-semantic dictionary construction as $\operatorname{KeySemanDictionary\_Constructor}()$ in Alg. 1. Although such a dictionary allows the subsequent attention operation to focus on the most semantically related patches, constructing it before each attention operation significantly increases computational costs. Meanwhile, we observed that IR architectures typically use transformers stage-by-stage (See Fig. 2(a)). This means that in each stage, several transformer layers are directly connected sequentially and operate at the same semantic level. Inspired by this, we propose to share the same key-semantic dictionary for all the transformer layers.
Algorithm 1 Key-Semantic Transformer Stage (i.e., SemanIR Stage)
0:  input feature ${F_{in}}$, numbers of SemanIR layer $N_{layer}$, KNN value $k$, the patched feature $\mathcal{P}$
0:  aggregated feature $F_{out}$
1:  $\mathcal{D}_{K}\leftarrow\operatorname{KeySemanDictionary\_Constructor}({\mathcal{P}},k)$ // Sec. 3.1
2:  for $i=1$ to $N_{layer}$  do
3:     $Q,K,V\leftarrow\operatorname{Linear\_Proj}({\mathcal{P}})$
4:     $\hat{{\mathcal{P}}}\leftarrow\operatorname{SemanIR\_Att}(Q,K,{V},\mathcal{D}_{K})$ // Sec. 3.2
5:     ${\mathcal{P}}\leftarrow\hat{{\mathcal{P}}}+\operatorname{FFN}(\hat{{\mathcal{P}}})$ // Sec. 3.2
6:  end for
7:  $F_{out}\leftarrow F_{in}+\operatorname{Conv}({\operatorname{UnPatch}(\mathcal{P})})$
8:  return $F_{out}$
3.2 Sharing Key Semantics Cross Transformer Layers
The structure of each transformer layer is shown in Fig. 2(b), which consists of a key-semantic attention block followed by a feed-forward network (FFN). Specifically, given an input $F_{in}$, we form each transformer layer as $z=\operatorname{FFN}(f_{\theta}(F_{in}))$, where $f$ means the transformer layer, $z$ denotes the output, and $\theta$ is the trainable parameters. Previous methods tried to reduce the computation cost mainly by applying some techniques (i.e., $\mathcal{T}$) like token merging or pruning after each attention calculation or the entire transformer layer, and it can be formalized as $z=\mathcal{T}(\operatorname{FFN}(f_{\theta}(F_{in})))$ or $z=(\operatorname{FFN}(\mathcal{T}f_{\theta}(F_{in})))$. However, the main computation cost from MSA is still untouched.
Owing to the permutation-invariant property (i.e., $f_{\theta}(\mathcal{T}x)=\mathcal{T}f_{\theta}(x)$, here $\mathcal{T}\in\mathbb{R}^{N\times N}$ means any token level permutation matrix) inherent in both the MSA and the FFN [89, 46], the transformer layer consistently produces identical representations for patches that share the same attributes, regardless of their positions or the surrounding structures [9]. In other words, patches at the same location are consistently connected to other patches possessing the same attributes as they traverse through the various layers within the same stage. It enables $\mathcal{D}_{K}$ to serve as a reference permutation for each attention in the subsequent transformer layers, facilitating efficient yet highly semantics-related attention operations. This distinguishes our method from previous token merging/pruning [110, 93] or sparse attention-based methods (Fig.1(d)) that only activate patches in a grid-fixed manner[102].
The workflow is intuitively illustrated in Fig. 2 (c) and (d). Initially, the patch $\mathcal{P}$ is linear projected via $\operatorname{Linear\_Proj()}$ (The 3rd step in Alg. 1) into $Q$, $K$, and $V$. For each patch $p_{i}$ in $Q$, instead of calculating the self-attention with all $hw$ patches in $K$&$V$, only $k$ essential patches are selected via the semantic lookup via the indices from $\mathcal{D}_{K}$ in them, forming the $\hat{K}$&$\hat{V}$. Then the attention matrix $\mathcal{D}_{K}^{att}$ is obtained by: $\mathcal{D}_{K}^{att}=\operatorname{Softmax_{K}}(Q\hat{K}^{\top}/\sqrt{d})$, which captures the pair-wise relation between each destination patch $p_{i}$ in $Q$ with only $k$ patches in $K$&$V$ that are semantically highly related to $p_{i}$. For other unselected patches in $K$&$V$, we aim to maintain their position in their corresponding places without any computation. Based on $\mathcal{D}_{K}^{att}$, the attention outputs the updated feature $\hat{\mathcal{P}}$ via: $\hat{\mathcal{P}}=\mathcal{D}_{K}^{att}\hat{V}$. We formulate these two procedures as $\operatorname{SemanIR\_Att()}$ in the 4th step of Alg. 1. This differs from the conventional MSA, which calculates the relation of each patch in $Q$ and all patches in $K$&$V$. Finally, with FFN, the output of each transformer layer is achieved via the 5th step of Alg. 1.
Conversely, our design offers two advantages. Firstly, the computational cost can be significantly reduced within each attention window (detailed analysis can be found in the Appx. A.3), enhancing efficiency. Additionally, sharing the key semantics across transformer layers within each stage acts as a loop that continuously optimizes a degraded patch with its most semantically related patches, ensuring the performance of the proposed method (supported by our experimental results in Sec. 4).
3.3 Discussion
Fixed top-k vs. Random top-k Training Strategies. In the fixed top-k approach, $k$ remains constant at 512 during training. In contrast, in the random top-k method, $k$ is randomly selected from the set $[64,128,192,256,384,512]$. It is important to note that even in the random top-k setting, a fixed k value is maintained for all patches/pixels in each iteration. During inference, $k$ is adjusted based on the computational constraints of users.
Implementation of the Attention of SemanIR. To achieve the attention operation of the proposed SemanIR, we explored three different manners for the implementation, i.e., (i) Triton, (ii) Torch-Gather, and (iii) Torch-Mask. Specifically, (i) is based on FlashAttention [20], and a customized GPU kernel is written for the operators proposed in this paper. Parallel GPU kernels are called for the nodes during run time. (ii) means that we use the ‘torch.gather()’ function in PyTorch to choose the corresponding $Q_{gather}$ and $K_{gather}$ based on $\mathcal{D}_{K}$, then the attention operation is conducted between $Q_{gather}$ and $K_{gather}$. (iii) denotes that we keep only the value of selected patches of $\mathcal{D}_{K}$ and omitting other patches with low correlation via assigning those values to $\infty$- ∞ guided by $\mathcal{D}_{K}$. Discussions of the pros and cons regarding these manners are provided in Sec. 4.1.
4 Experiments
Figure 3: The impact of $k$ with different inference $k$ value. Circle size represents FLOPs.
Table 1: GPU memory footprint of different implementations (i.e., Triton, Torch-Gather, and Torch-Mask) of our key-graph attention block. $N$ is the number of tokens and $k$ is the number of nearest neighbors. OOM denotes "out of memory".
$N$ Triton Torch-Gather Torch-Mask
512 0.27 GB 0.66 GB 0.36 GB
1024 0.33 GB 1.10 GB 0.67 GB
2048 0.68 GB 2.08 GB 1.91 GB
4096 2.61 GB 4.41 GB 6.83 GB
8192 10.21 GB 10.57 GB 26.42 GB
$k$ Triton Torch-Gather Torch-Mask
32 5.51 GB 15.00 GB 13.68 GB
64 5.82 GB 27.56 GB 13.93 GB
128 6.45 GB OOM 14.43 GB
256 7.70 GB OOM 15.43 GB
512 10.20 GB OOM 17.43 GB
Table 2: Single-image motion deblurring results. GoPro [67] dataset is used for training.
GoPro HIDE Average
[…]
DeblurGAN [41] 28.70 0.858 24.51 0.871 26.61 0.865
Nah et al.  [67] 29.08 0.914 25.73 0.874 27.41 0.894
DeblurGAN-v2 [42] 29.55 0.934 26.61 0.875 28.08 0.905
SRN [85] 30.26 0.934 28.36 0.915 29.31 0.925
Gao et al. [25] 30.90 0.935 29.11 0.913 30.01 0.924
DBGAN [107] 31.10 0.942 28.94 0.915 30.02 0.929
MT-RNN [69] 31.15 0.945 29.15 0.918 30.15 0.932
DMPHN [100] 31.20 0.940 29.09 0.924 30.15 0.932
Suin et al. [84] 31.85 0.948 29.98 0.930 30.92 0.939
CODE [113] 31.94 - 29.67 - 30.81 -
SPAIR [70] 32.06 0.953 30.29 0.931 31.18 0.942
MIMO-UNet+ [17] 32.45 0.957 29.99 0.930 31.22 0.944
IPT [10] 32.52 - - - - -
MPRNet [98] 32.66 0.959 30.96 0.939 31.81 0.949
KiT [45] 32.70 0.959 30.98 0.942 31.84 0.951
Restormer [97] 32.92 0.961 31.22 0.942 32.07 0.952
Ren et al.  [76] 33.20 0.963 30.96 0.938 32.08 0.951
SemanIR (ours) 33.44 0.964 31.05 0.941 32.25 0.953
Table 3: The efficiency comparisons results on Urban100 dataset.
[…]
×$\times 4$ SR SwinIR [52] Columnar 11.90 215.32 152.24 27.45
CAT [16] Columnar 16.60 387.86 357.97 27.89
HAT [14] Columnar 20.77 416.90 368.61 28.37
SemanIR-S (Ours) Columnar 12.02 290.20 211.94 28.34
Denoising ($\sigma=50$) (The same architecture for other IR task) SwinIR [52] Columnar 11.75 752.06 1772.84 27.98
Restormer [97] U-shape 26.10 154.88 210.44 28.29
GRL [50] Columnar 19.81 1361.77 3944.17 28.59
SemanIR (Ours) U-shape 25.85 135.26 240.05 28.63
In this section, we first analyze three important ablation studies of our SemanIR, followed by extensive experiments on 6 IR tasks, i.e., deblurring, JPEG CAR, image denoising, IR in AWC, image demosaicking, and image SR. More details about the architecture design, training protocols, the training/testing dataset, and full quantitative/additional qualitative results are shown in Appx. A to  E. The best and the 2nd-best results are reported in red and blue, respectively. Note that denotes a single model that is trained to handle multiple degradation levels i.e., noise levels, and quality factors.
4.1 Ablation Study
The impact of the implementation of Key-Graph Attention is assessed in terms of (i) Triton, (ii) Torch-Gather, and (iii) Torch-Mask under different numbers of N (various from 512 to 8192) and K (various from 32 to 512). The results of the GPU memory footprint are shown in Tab. 2, which indicate that Torch-Gather brings no redundant computation while requiring a large memory footprint. Though Torch-Mask brings the GPU memory increase, the increment is affordable compared to Torch-Gather and also easy to implement. Triton largely saves the GPU memory while at the cost of slow inference and difficult implementation for the back-propagation process. To optimize the efficiency of our SemanIR, we recommend employing Torch-Mask during training and Triton during inference, striking a balance between the efficiency and the GPU memory requirement.
The Impact of the $k$ in Key-Semantic Dictionary Construction. Three interesting phenomena are observed from the results shown in Fig. 3 regarding the two top-k training strategies (Sec. 3.3). (1) The PSNR can largely increase with the increase of $k$ in a fixed manner. (2) When $k$ reaches a certain number (i.e., 384), the performance improvements become marginal, supporting our statement that only the most semantically related patches contribute significantly to the restoration. (3) The randomly sampled strategy has a very stable and better performance compared to the fixed top-k manner especially when the inference $k$ is fixed to a small number (i.e., 64, 128, 256). We conclude that a random sampled strategy is more general and stable. It can also make the inference process more flexible regarding different computation resources. More ablation results can be found in our Appx. D about the effect of the noise level and quality factor for denoising and JPEG CAR.
Table 4: Color image JPEG compression artifact removal results.
Set QF JPEG QGAC [22] FBCNN [35] DRUNet [103] SemanIR (Ours) SwinIR [52] GRL-S [50] SemanIR (Ours)
[…]
LIVE1 10 25.69 0.7430 27.62 0.8040 27.77 0.8030 27.47 0.8045 28.19 0.8146 28.06 0.8129 28.13 0.8139 28.31 0.8176
20 28.06 0.8260 29.88 0.8680 30.11 0.8680 30.29 0.8743 30.53 0.8781 30.44 0.8768 30.49 0.8776 30.61 0.8792
30 29.37 0.8610 31.17 0.8960 31.43 0.8970 31.64 0.9020 31.89 0.9051 31.81 0.9040 31.85 0.9045 31.94 0.9058
40 30.28 0.8820 32.05 0.9120 32.34 0.9130 32.56 0.9174 32.81 0.9201 32.75 0.9193 32.79 0.9195 32.85 0.9204
BSDS500 10 25.84 0.7410 27.74 0.8020 27.85 0.7990 27.62 0.8001 28.25 0.8076 28.22 0.8075 28.26 0.8083 28.37 0.8102
20 28.21 0.8270 30.01 0.8690 30.14 0.8670 30.39 0.8711 30.55 0.8738 30.54 0.8739 30.57 0.8746 30.63 0.8750
30 29.57 0.8650 31.33 0.8980 31.45 0.8970 31.73 0.9003 31.90 0.9026 31.90 0.9025 31.92 0.9030 31.96 0.9035
40 30.52 0.8870 32.25 0.9150 32.36 0.9130 32.66 0.9168 32.84 0.9190 32.84 0.9189 32.86 0.9192 32.88 0.9193
Table 5: IR in adverse weather conditions.
Type Test1 (rain+fog) SnowTest100k-L RainDrop
[…]
Task Specific pix2pix [33] 19.09 DesnowNet [56] 27.17 AttGAN [71] 30.55
HRGAN [47] 21.56 JSTASR [12] 25.32 Quan [73] 31.44
SwinIR [52] 23.23 SwinIR 28.18 SwinIR 30.82
MPRNet [98] 21.90 DDMSNET [106] 28.85 CCN [72] 31.34
Multi Task All-in-One [48] 24.71 All-in-One 28.33 All-in-One 31.12
TransWea. [88] 27.96 TransWea. 28.48 TransWea. 28.84
SemanIR (Ours) 29.57 SemanIR (Ours) 30.76 SemanIR (Ours) 30.82
Table 6: Image demosaicking PSNR results.
Datasets Kodak McMaster
Matlab 35.78 34.43
MMNet [40] 40.19 37.09
DDR [92] 41.11 37.12
DeepJoint [27] 42.00 39.14
RLDD [30] 42.49 39.25
DRUNet [103] 42.68 39.39
RNAN [112] 43.16 39.70
GRL [50] 43.57 40.22
SemanIR (Ours) 43.62 40.68
Efficiency Analysis. We compare our SemanIR with 4 recent promising methods both the x4 SR and grayscale denoising on the Urban100. The trainable parameters, the FLOPs, the runtime, and the PSNR are reported in Tab. 3. It shows that: (1) Among all the methods, HAT and SemanIR achieve 1st-class PSNR performances, reaching 28.37dB and 28.34dB, while SemanIR-S is much faster and has 41.7% fewer parameters than HAT. (2) SwinIR runs a bit faster than SemanIR-S but with a PSNR loss of 0.89 dB. Compared with CAT, and HAT, our SemanIR-S is faster and more accurate. In addition, the parameters and the FLOPs comparison between SwinIR and our SemanIR for denoising reveals that despite SwinIR having fewer training parameters, its FLOPs significantly surpass ours.
Figure 4: One model is trained to handle multi-degradations for denoising (a-b) and JPEG CAR (c-d).
The Impact of One Model is Trained to Handle Multiple Degradation Levels. The experiments were conducted for both denoising and JPEG CAR with both color and grayscale images. For denoising, $\sigma$ is set to $[15,25,50,75]$. For JEPG CAR, QF is set to $[10,20,30,40,50,60,70,80,90]$. The results in Fig. 4 indicate that the PSNR for both tasks across all the datasets, under both color and grayscale settings, decreases when the degraded level increases. It means to maintain a decent generalization ability with one model to handle multiple degradation levels is not easy. Meanwhile, the proposed SemanIR can still outperform other methods on various tasks (See Tab. 7 and Tab. 4), which means that capturing the key semantics is also essential for model’s generalization ability.
4.2 Evaluation of SemanIR on Various IR Tasks
Evaluation on Image deblurring. Tab. 2 shows the quantitative results for single image motion deblurring on synthetic datasets (GoPro [67], HIDE [81]). Compared to the previous state-of-the-art Restormer [97], our SemanIR achieves significant PSNR improvement (i.e., 0.52 dB) on the GoPro dataset and the second-best on the HIDE dataset. Visual results are shown in the Appx. E.2.
Table 7: Color and grayscale image denoising PSNR results.
Method # P Color Grayscale
CBSD68 McMaster Urban100 Set12 BSD68 Urban100
$\sigma$$=$$15$ $\sigma$$=$$25$ $\sigma$$=$$50$ $\sigma$$=$$15$ $\sigma$$=$$25$ $\sigma$$=$$50$ $\sigma$$=$$15$ $\sigma$$=$$25$ $\sigma$$=$$50$ $\sigma$$=$$15$ $\sigma$$=$$25$ $\sigma$$=$$50$ $\sigma$$=$$15$ $\sigma$$=$$25$ $\sigma$$=$$50$ $\sigma$$=$$15$ $\sigma$$=$$25$ $\sigma$$=$$50$
DnCNN [38] 0.56 33.90 31.24 27.95 33.45 31.52 28.62 32.98 30.81 27.59 32.67 30.35 27.18 31.62 29.16 26.23 32.28 29.80 26.35
FFDNet [105] 0.49 33.87 31.21 27.96 34.66 32.35 29.18 33.83 31.40 28.05 32.75 30.43 27.32 31.63 29.19 26.29 32.40 29.90 26.50
DRUNet [103] 32.64 34.30 31.69 28.51 35.40 33.14 30.08 34.81 32.60 29.61 33.25 30.94 27.90 31.91 29.48 26.59 33.44 31.11 27.96
Restormer [97] 26.13 34.39 31.78 28.59 35.55 33.31 30.29 35.06 32.91 30.02 33.35 31.04 28.01 31.95 29.51 26.62 33.67 31.39 28.33
SemanIR (Ours) 25.82 34.42 31.78 28.57 35.65 33.40 30.34 35.37 33.26 30.41 33.47 31.16 28.12 31.95 29.49 26.54 34.05 31.84 28.83
DnCNN [38] 0.56 33.90 31.24 27.95 33.45 31.52 28.62 32.98 30.81 27.59 32.86 30.44 27.18 31.73 29.23 26.23 32.64 29.95 26.26
EDT-B [49] 11.48 34.39 31.76 28.56 35.61 33.34 30.25 35.22 33.07 30.16 - - - - - - - - -
DRUNet [103] 32.64 34.30 31.69 28.51 35.40 33.14 30.08 34.81 32.60 29.61 33.25 30.94 27.90 31.91 29.48 26.59 33.44 31.11 27.96
SwinIR [52] 11.75 34.42 31.78 28.56 35.61 33.20 30.22 35.13 32.90 29.82 33.36 31.01 27.91 31.97 29.50 26.58 33.70 31.30 27.98
Restormer [97] 26.13 34.40 31.79 28.60 35.61 33.34 30.30 35.13 32.96 30.02 33.42 31.08 28.00 31.96 29.52 26.62 33.79 31.46 28.29
Xformer [101] 25.23 34.43 31.82 28.63 35.68 33.44 30.38 35.29 33.21 30.36 33.46 31.16 28.10 31.98 29.55 26.65 33.98 31.78 28.71
SemanIR (Ours) 25.82 34.43 31.79 28.60 35.65 33.43 30.38 35.38 33.29 30.51 33.48 31.18 28.14 31.97 29.52 26.53 34.09 31.87 28.86
Evaluation on JPEG CAR. The experiments for both grayscale and color images are conducted with 4 image quality factors ranging from 10 to 40 under two settings (i.e., a single model is trained to handle multiple quality factors, and each model for each quality). The quantitative results shown in Tab. 4 and the Appx. E.1 show that our SemanIR achieves the best results on all the test sets cross various quality factors among all the comparison methods for both the color and grayscale images. The visual comparisons in the Appx. E.2 further supports the effectiveness of our method.
Evaluation on Image Denoising. We show color and grayscale image denoising results in Tab. 7 under two settings (i.e., one model for all noise levels $\sigma=\{15,25,50\}$ and each model for each noise level). For a fair comparison, both parameters and accuracy are reported for all the methods. For , our SemanIR performs better on all test sets for color and grayscale image denoising than others. It is worth noting that we outperform DRUNet and Restormer with lower trainable parameters. For another setting, the proposed SemanIR also archives better results on CBSD68 and Urban100 for color image denoising, and on Set12 and Urban100 for grayscale denoising. These interesting comparisons validate the effectiveness of the proposed SemanIR and also indicate that our method has a higher generalization ability. The visual results in Appx. E.2 also support that the proposed SemanIR can remove heavy noise corruption and preserve high-frequency image details, resulting in sharper edges and more natural textures without over-smoothness or over-sharpness problems.
Table 8: Classical image SR results. Both lightweight and accurate models are summarized.
Method Scale Params Set5 Set14 BSD100 Urban100 Manga109
[…]
RCAN [111] 2$\times$ 15.44 38.27 0.9614 34.12 0.9216 32.41 0.9027 33.34 0.9384 39.44 0.9786
SAN [19] 2$\times$ 15.71 38.31 0.9620 34.07 0.9213 32.42 0.9028 33.10 0.9370 39.32 0.9792
HAN [68] 2$\times$ 63.61 38.27 0.9614 34.16 0.9217 32.41 0.9027 33.35 0.9385 39.46 0.9785
IPT [10] 2$\times$ 115.48 38.37 - 34.43 - 32.48 - 33.76 - - -
SwinIR [52] 2$\times$ 11.75 38.42 0.9623 34.46 0.9250 32.53 0.9041 33.81 0.9427 39.92 0.9797
CAT-A [16] 2$\times$ 16.46 38.51 0.9626 34.78 0.9265 32.59 0.9047 34.26 0.9440 40.10 0.9805
ART [102] 2$\times$ 16.40 38.56 0.9629 34.59 0.9267 32.58 0.9048 34.30 0.9452 40.24 0.9808
EDT [49] 2$\times$ 11.48 38.63 0.9632 34.80 0.9273 32.62 0.9052 34.27 0.9456 40.37 0.9811
SemanIR-S (Ours) 2$\times$ 11.87 38.57 0.9651 34.99 0.9300 32.65 0.9078 34.86 0.9472 40.45 0.9824
SemanIR-B (Ours) 2$\times$ 19.90 38.61 0.9654 35.08 0.9304 32.69 0.9084 34.99 0.9455 40.59 0.9830
RCAN [111] 4$\times$ 15.59 32.63 0.9002 28.87 0.7889 27.77 0.7436 26.82 0.8087 31.22 0.9173
SAN [19] 4$\times$ 15.86 32.64 0.9003 28.92 0.7888 27.78 0.7436 26.79 0.8068 31.18 0.9169
HAN [68] 4$\times$ 64.20 32.64 0.9002 28.90 0.7890 27.80 0.7442 26.85 0.8094 31.42 0.9177
IPT [10] 4$\times$ 115.63 32.64 - 29.01 - 27.82 - 27.26 - - -
SwinIR [52] 4$\times$ 11.90 32.92 0.9044 29.09 0.7950 27.92 0.7489 27.45 0.8254 32.03 0.9260
CAT-A [16] 4$\times$ 16.60 33.08 0.9052 29.18 0.7960 27.99 0.7510 27.89 0.8339 32.39 0.9285
ART [102] 4$\times$ 16.55 33.04 0.9051 29.16 0.7958 27.97 0.751 27.77 0.8321 32.31 0.9283
EDT [49] 4$\times$ 11.63 33.06 0.9055 29.23 0.7971 27.99 0.7510 27.75 0.8317 32.39 0.9283
SemanIR-S (Ours) 4$\times$ 12.02 33.02 0.9082 29.29 0.8026 27.96 0.7582 28.34 0.8467 32.48 0.9322
SemanIR-B (Ours) 4$\times$ 20.04 33.08 0.9090 29.34 0.8037 27.98 0.7599 28.51 0.8467 32.56 0.9335
Figure 5: Visual comparison of classical image SR (4$\times$) on Urban100. Best viewed by zooming.
Evaluation in AWC. We validate SemanIR in adverse weather conditions, including rain+fog (Test1), snow (SnowTest100K), and raindrops (RainDrop). PSNR is reported in Tab. 6. Our method achieves the best performance on Test1 (i.e., 5.76% improvement) and SnowTest100k-L (i.e. 8.01% improvement), while the second-best PSNR on RainDrop compared to all other methods. See Appx. E.2 for Visual comparisons.
Evaluation on Image Demosaicking. The quantitative results shown in 6 indicate that the proposed SemanIR performs best on both the Kodak and MaMaster test sets, especially 0.05dB and 0.45dB absolute improvement compared to the current state-of-the-art.
Evaluation on SR. For the classical image SR, we compared our SemanIR with both recent lightweight and accurate SR models, and the quantitative results are shown in Tab. 8. Compared to EDT, SemanIR-base achieves significant improvements on Urban100 (i.e., 0.72 dB and 0.76dB for 2$\times$ and 4$\times$ SR) and Manga109 datasets (i.e., 0.22dB and 0.17 dB for 2$\times$ and 4$\times$ SR). Even the SemanIR-small consistently ranks as the runner-up across the majority of test datasets, all while maintaining a reduced number of trainable parameters. Visual results in Fig. 5 and Appx. E.2 also validate the effectiveness of the proposed SemanIR in restoring more details and structural content.
5 Conclusion
In this paper, we propose a novel approach, SemanIR, for ViTs-based image restoration, which experimentally validated that global cues are essential to restore degraded images well, but the most semantically related global cures play the major role. Specifically, to capture the key semantics, we propose to construct a semantic dictionary (i.e., naively by self-similarity is enough) for storing only the most related $k$ semantic information and then use it as a reference for guiding the attention operation for making the attention operation pay more attention only to these key semantics. Furthermore, we share the key-semantic dictionary with all the upcoming transformer layers within the same stage since each stage of the transformer is typically at the same semantic level. This strategy significantly reduces the computational cost for IR and functions as loop optimization, continuously restoring degraded patches with their most semantically related patches, which share similar texture or structural information. Extensive experiments on 6 IR tasks validated the effectiveness of SemanIR, demonstrating that our method achieves new state-of-the-art performance.
Appendix A Experimental Protocals
A.1 Training/Testing Datasets
JPEG compression artifact removal. For JPEG compression artifact removal, the JPEG image is compressed by the cv2 JPEG compression function. The compression function is characterized by the quality factor. We investigated four compression quality factors including 10, 20, 30, and 40. The smaller the quality factor, the more the image is compressed, meaning a lower quality.
  • The training datasets: DIV2K [1], Flickr2K [53], and WED [61].
  • The test datasets: Classic5 [23], LIVE1 [80], Urban100 [32], BSD500 [2].
Image Denoising. For image denoising, we conduct experiments on both color and grayscale image denoising. During training and testing, noisy images are generated by adding independent additive white Gaussian noise (AWGN) to the original images. The noise levels are set to $\sigma=15,25,50$. We train individual networks at different noise levels. The network takes the noisy images as input and tries to predict noise-free images.
  • The training datasets: DIV2K [1], Flickr2K [53], WED [61], and BSD400 [62].
  • The test datasets for color image: CBSD68 [62], Kodak24 [24], McMaster [109], and Urban100 [32].
  • The test datasets for grayscale image: Set12 [104], BSD68 [62], and Urban100 [32].
Image Demosaicking. For image demosaicking, the mosaic image is generated by applying a Bayer filter on the ground-truth image. Then the network try to restore high-quality image. The mosaic image is first processed by the default Matlab demosaic function and then passed to the network as input.
  • The training datasets: DIV2K [1] and Flickr2K [53].
  • The test datasets: Kodak [24], McMaster [109].
IR in Adverse Weather Conditions. For IR in adverse weather conditions, the model is trained on a combination of images degraded by a variety of adverse weather conditions. The same training and test dataset is used as in Transweather [88]. The training data comprises 9,000 images sampled from Snow100K [56], 1,069 images from Raindrop [71], and 9,000 images from Outdoor-Rain [47]. Snow100K includes synthetic images degraded by snow, Raindrop consists of real raindrop images, and Outdoor-Rain contains synthetic images degraded by both fog and rain streaks. The proposed method is tested on both synthetic and real-world datasets.
  • The comparison methods in Tab. 6 of our main manuscript: pix2pix [33], HRGAN [47], SwinIR [52], All-in-One [48], Transweather [88], DesnowNet [56], JSTASR [12], DDMSNET [106], Attn. GAN [71][73], and CCGAN [72].
  • The test datasets: test1 dataset [48, 47], the RainDrop test dataset [71], and the Snow100k-L test.
Image SR. For image SR, the LR image is synthesized by Matlab bicubic downsampling function before the training. We investigated the upscalingg factors ×$\times 2$, ×$\times 3$, and ×$\times 4$.
  • The training datasets: DIV2K [1] and Flickr2K [53].
  • The test datasets: Set5 [4], Set14 [99], BSD100 [62], Urban100 [32], and Manga109 [63].
Image Deblurring. For single-image motion deblurring,
  • The training datasets: GoPro [67] .
  • The test datasets: GoPro [67] and HIDE [81].
Figure 6: The U-shaped hierarchical architecture (Archi-V2) of the proposed SemanIR for Image Restoration. Note that this U-shaped one is used for image JPEG CAR, image denoising, image demosaicking, IR in AWC, and image deblurring. Symbols $\oplus$ and ⓒ denote the element-wise addition and channel-wise concatenation. The downsample and upsample operations are denoted by red and green arrows.
Table 9: The details of the transformer stages and layers per stage of SemanIR for both architectures.
Archi-V1 (Columnar-shape) Archi-V2 (U-shape)
SemanIR-small SemanIR-base Down Stages Up Stages Final Stage
Num. of SemanIR Stages 6 8 4 4 1
Num. of SemanIR layer per stage 6 8 6 6 6
A.2 Model Architecture
In the proposed SemanIR, we adopt two kinds of base architecture i.e., the widely used multi-stage one shown in Fig.1 of our main manuscript (Archi-V1) and a U-shaped hierarchical one shown in Fig. 6 (Archi-V2) for taking patterns of various scales into account (Note that 1/3 of $I_{low}$ and $I_{high}$ in Fig. 6 denotes the grayscale/color image cases). This is consistent with previous methods such as Restormer [97], KiT [45], and NAFNet [11].
Note that The feature extractor for both architectures is implemented as a simple convolution and converts the input image into feature maps. The image reconstructor for Archi-V1 takes the rich features calculated by the previous operations and estimates a recovered image.
In addition to introducing the two base architectures of the proposed SemanIR, we have provided comprehensive details of its structure in Table 9. This table outlines the number of SemanIR stages and the distribution of layers within each stage, offering a thorough understanding of our model’s architecture.
A.3 Efficiency Analysis
Table 10: The Computation Complexity Comparison.
Time Complexity Space Complexity
MSA $\mathcal{O}(4HWC^{2}+2(HW)^{2}C)$ $\mathcal{O}(4HWC^{2}+2h(HW)^{2}C)$
W-MSA $\mathcal{O}(4HWC^{2}+2(M)^{2}HWC)$ $\mathcal{O}(4HWC^{2}+2h(M)^{2}HWC)$
SemanIR MSA (Ours) $\mathcal{O}(4HWC^{2}+2kHWC)$ $\mathcal{O}(4HWC^{2}+2hkHWC)$
We provide a complexity comparison among the standard multi-head self-attention (MSA), the Window-wise MSA (W-MSA), and the proposed KeySemanIR MSA (SemanIR-MSA) in the Tab. 10. $(H,W,C)$ indicate the feature size, $M$ represents the window size, and $h$ denotes the number of heads. It is commonly demonstrated and proven that the complexity of the W-MSA is much lower than that of the standard MSA, i.e.,
$\mathcal{O}(4HWC^{2}+2(M)^{2}HWC)<\mathcal{O}(4HWC^{2}+2(HW)^{2}C)$ (3)
To better understand the efficiency of the proposed method, it should be considered together with all transformer layers within a certain stage. Specifically, take one stage, which contains 6 transformer Layers, as an example (To simplify the illustration, we omit the convolution operation at the end of each stage).
First, the total complexity of W-MSA within each stage can be calculated as:
$\mathcal{O}(6\times[4HWC^{2}+2(M)^{2}HWC])$ (4)
Second, similarly, the complexity of the proposed SemanIR-MSA can be calculated as follows. ($\mathcal{O}(HWC)$ indicates the complexity of the key-semantic dictionary construction at the start of each transformer stage. All the other layers then share it, hence it is calculated only once).
$\mathcal{O}(6\times[4HWC^{2}+2kHWC]+HWC)$ (5)
Third, a simple subtraction can be done as follows to validate that the proposed method is more efficient compared to the W-MSA within each Transformer stage:
$\mathcal{O}(6\times[4HWC^{2}+2kHWC]+HWC-6\times[4HWC^{2}+2(M)^{2}HWC])$ (6)
$=\mathcal{O}((12M^{2}-K-1)HWC)>>0$
In the last equation, since $M^{2}$ contains all the patches/pixels within a certain window, for example, in a common setting with a window size of M = 7 and a patch size of 16, there are around (7 $\times$ 7) $\times$ (16 $\times$ 16) pixels within the given window. However, the k value in our proposed Key-Semantic dictionary is set to 512 or randomly sampled from [64, 128, 256, 384, 512], which is significantly smaller than (7 $\times$ 7) $\times$ (16 $\times$ 16).
Based on the above analysis, it can be concluded that, together with the proposed transformer layer, constructing the key-semantic dictionary at the start of each stage leads to greater efficiency.
A.4 Training Details
Our method explores 6 IR tasks, and the training settings vary slightly for each task. These differences encompass the architecture of the proposed SemanIR, variations in the choice of the optimizer, and loss functions.
Figure 7: Ablation study on the impact of $k$ for Color JPEG CAR on CBSD68 (a-d), Live (e-h), and Urban100 (i-l) datasets with $QF$ = {10, 20, 30, 40}.
Architecture. We use the columnar multi-stage architecture (without changing the feature map resolution and number of channels) for image SR and the U-shaped architecture for the other tasks including image denoising, image deblurring, and other tasks. The strategy of using multiple architectures is also explored by the previous method [14, 11].
Optimizer. We adopt the same optimizer as all other comparison methods, i.e., Adam [39], for IR in AWC, and AdamW [59] for the rest IR tasks.
Loss Function. We adopt the same loss function as all other comparison methods, i.e., smooth L1 loss and VGG loss [36, 83] for IR in AWC, the Charbonnier loss for Deblurring, and L1 loss for the rest IR tasks.
Figure 8: Ablation study on the impact of $k$ for Grayscale JPEG CAR on BSD500 (a-d), Classic5 (e-h), Live1(i-l), and Urban100 (m-p) datasets with $QF$ = {10, 20, 30, 40}.
Batch Size and Patch Size. We keep the similar batch size as other comparison methods, i.e., (Batch size = 16, Patch Size = 64) for JPEG CAR, denoising, demosaicking, and SR. (Batch Size = 32, Patch Size = 16) for IR in AWC. (Batch Size = 8, Patch Size = 192) for deblurring.
Learning Rate Schedule. For all the IR tasks, similar to other comparison methods, we set the initial learning rate to 2 $\times$ 10-4, and then the half-decay is adopted during the training. Note that the training iteration for JPEG CAR, denoising, demosaicking, and SR is set to 1M. For IR in AWC and debluriing, it is set to 750K.
A.5 Evaluation Introduction
Note that the results of all the comparison methods are reported from their original papers. The details of the evaluation metric (i.e., SSIM, PSNR) are described as follows:
JPEG compression artifact removal. For color image JPEG compression artifact removal, the PSNR is reported on the RGB channels while for grayscale image JPEG compression artifact removal, the PSNR is reported on the Y channel.
Image Denoising. For color image denoising, the PSNR is reported on the RGB channels while for grayscale image denoising, the PSNR is reported on the Y channel.
Image Demosaicking. For the comparison between different methods, PSNR is reported on the RGB channels.
IR in Adverse Weather Conditions. We adopted the same PSNR evaluation metric used in Transweather [88].
Image SR. The PSNR is reported on the Y channel.
Image Deblurring. The PSNR and SSIM on the RGB channels are reported.
Figure 9: Ablation study on the impact of $k$ for Color Image Denoising on CBSD68 (a-c), Kodak24 (d-f), McMaster (g-i), and Urban100 (j-l) datasets with $\sigma$ = {15, 25, 50}.
Figure 10: Training log shows the convergence of the proposed SemanIR during training. The upscaling factor is 2$\times$.
Appendix B Limitations
This study faces a task-specific limitation: each image restoration task requires training a separate network. While efforts have been made to train models for varying degradation levels within specific types, such as image denoising and removal of JPEG compression artifacts, this approach still leads to inefficiencies in model training. It constrains the utility of the trained networks. A potential future enhancement involves developing a mechanism enabling a network to handle diverse image degradation types and levels. Another challenge is the substantial parameter requirement of the proposed SemanIR, which operates within a tens-of-millions parameter budget. Deploying such a large image IR network on handheld devices with limited resources is challenging, if not unfeasible. Therefore, a promising research direction is the creation of more efficient versions of SemanIR, integrating non-local context more effectively, to overcome these limitations.
Appendix C Impact Statement
This paper introduces a transformer-based approach that significantly enhances the efficiency and performance of image restoration tasks, including image deblurring, JPEG CAR, image denoising, IR in adverse weather conditions, demosaicking, and image super-resolution. The proposed SemanIR’s notable efficiency improvement holds promise for resource-effective implementations in real-world applications. This elevated performance creates opportunities for enhanced image quality across diverse domains. While our primary contribution lies in the technical aspects of Machine Learning, we are cognizant of potential societal impacts, particularly in healthcare, surveillance, and digital imaging. As with any technology, ongoing vigilance in ethical considerations during deployment is essential, ensuring responsible use and proactively addressing any unintended consequences.
Appendix D More Ablation Analyses
Besides the ablation studies presented in our main manuscript, we further provide the following two analyses:
Figure 11: Ablation study on the impact of $k$ for Grayscale Image Denoising on BSD68 (a-c), Set12 (d-f), and Urban100 (g-i) datasets with $\sigma$ = {15, 25, 50}.
Convergence Visualization. The training log of the proposed SemanIR for image SR is shown in Fig. 10. The log is reported for the PSNR on the Set5 dataset during training. Two versions of the proposed method including SemanIR-S and SemanIR-B are shown in this figure. As shown in this figure, the proposed network converges gradually during the training.
The Impact of the $k$ in Key-Semantic Dictionary Construction under Various IR Tasks. To explore how the $k$ value of top-k will affect the IR performance of the proposed SemanIR. We conduct exhaustive experiments on JPEG compression artifact reduction for both color and grayscale images under different QF values (i.e., QF = [10, 20, 30, 40]), image denoising for both color and grayscale images under different noise levels (i.e., $\sigma$ = [15, 25, 50]), as well as image SR under different scales (i.e., 2$\times$, 3$\times$, 4$\times$) with the proposed SemanIR. Note that all the experiments for each IR task are conducted under two kinds of top-k settings, i.e., (i) $k$ was randomly sampled from the range [64, 512] during the overall training phase, and (ii) $k$ was held constant at 512 throughout the training phase. For inference, $k$ was configured to the specified value for both settings.
The results of the JPEG CAR in terms of the hyper-parameters $k$ under different training settings during inference for both color image and grayscale image are shown in Fig. 7 and Fig. 8. It is clear that for both the color and grayscale JPEG CAR task, when $k$ is set to 64 during inference, there is a huge performance cat between the random top-k setting and the fixed top-k setting. In addition, the fixed top-k setting performs well or sometimes even a bit better than the random top-k setting only when $k$ is also set to the same number (i.e., 512). With the decrease of $k$ during inference for the fixed top-k setting, the PSNR drops largely marginally for all the datasets under every kind of degraded QF factor.
[…]
Table 11: Full Grayscale image JPEG compression artifact removal results.
Set QF JPEG DnCNN3 [104] DRUNet [103] SemanIR (Ours) GRL-S [50] SwinIR [52] ART [102] CAT [16] SemanIR (Ours)
[…]
Classic5 10 27.82 0.7600 29.40 0.8030 30.16 0.8234 30.26 0.8240 30.20 0.8286 30.27 0.8249 30.27 0.8258 30.26 0.8250 30.36 0.8267
20 30.12 0.8340 31.63 0.8610 32.39 0.8734 32.52 0.8740 32.49 0.8776 32.52 0.8748 - - 32.57 0.8754 32.58 0.8748
30 31.48 0.8670 32.91 0.8860 33.59 0.8949 33.74 0.8955 33.72 0.8985 33.73 0.8961 33.74 0.8964 33.77 0.8964 33.77 0.8958
40 32.43 0.8850 33.77 0.9000 34.41 0.9075 34.55 0.9078 34.53 0.9107 34.52 0.9082 34.55 0.9086 34.58 0.9087 34.57 0.9080
LIVE1 10 27.77 0.7730 29.19 0.8120 29.79 0.8278 29.84 0.8323 29.82 0.8323 29.86 0.8287 29.89 0.8300 29.89 0.8295 29.92 0.8360
20 30.07 0.8510 31.59 0.8800 32.17 0.8899 32.23 0.8949 32.22 0.8930 32.25 0.8909 - - 32.30 0.8913 32.28 0.8950
30 31.41 0.8850 32.98 0.9090 33.59 0.9166 33.65 0.9213 33.65 0.9190 33.69 0.9174 33.71 0.9178 33.73 0.9177 33.69 0.9201
40 32.35 0.9040 33.96 0.9250 34.58 0.9312 34.65 0.9329 34.64 0.9331 34.67 0.9317 34.70 0.9322 34.72 0.9320 34.67 0.9345
Urban100 10 26.33 0.7816 28.54 0.8484 30.31 0.8745 30.81 0.8885 30.70 0.8875 30.55 0.8835 30.87 0.8894 30.81 0.8866 31.15 0.8941
20 28.57 0.8545 31.01 0.9050 32.81 0.9241 33.33 0.9266 33.24 0.9270 33.12 0.9190 - - 33.38 0.9269 33.51 0.9272
30 30.00 0.9013 32.47 0.9312 34.23 0.9414 34.74 0.9446 34.67 0.9430 34.58 0.9417 34.81 0.9442 34.81 0.9449 34.84 0.9462
40 31.06 0.9215 33.49 0.9412 35.20 0.9547 35.69 0.9447 35.62 0.9519 35.50 0.9515 35.73 0.9553 35.73 0.9511 35.75 0.9550
The results of the image denoising in terms of hyper-parameters $k$ under different training settings during inference for both color image and grayscale image are shown in Fig. 9 and Fig. 11. All the experimental results on various datasets (i.e., BSD68/CBSD68, Kodak24, McMaster, and Urban100) share a similar trend for image denoising compared to the JPEG CAR task. The random top-k setting can maintain a relatively stable PSNR score under different $k$ during inference compared to its fixed counterpart. In addition, a decent result can be obtained for the fixed top-k setting only when the $k$ is set to the same (i.e., 512) during the inference.
Table 12: Full Classical image SR results. Both lightweight and accurate models are summarized.
Method Scale Params Set5 Set14 BSD100 Urban100 Manga109
[…]
RCAN [111] 2$\times$ 15.44 38.27 0.9614 34.12 0.9216 32.41 0.9027 33.34 0.9384 39.44 0.9786
SAN [19] 2$\times$ 15.71 38.31 0.9620 34.07 0.9213 32.42 0.9028 33.10 0.9370 39.32 0.9792
HAN [68] 2$\times$ 63.61 38.27 0.9614 34.16 0.9217 32.41 0.9027 33.35 0.9385 39.46 0.9785
NLSA [64] 2$\times$ 42.63 38.34 0.9618 34.08 0.9231 32.43 0.9027 33.42 0.9394 39.59 0.9789
IPT [10] 2$\times$ 115.48 38.37 - 34.43 - 32.48 - 33.76 - - -
SwinIR [52] 2$\times$ 11.75 38.42 0.9623 34.46 0.9250 32.53 0.9041 33.81 0.9427 39.92 0.9797
CAT-A [16] 2$\times$ 16.46 38.51 0.9626 34.78 0.9265 32.59 0.9047 34.26 0.9440 40.10 0.9805
ART [102] 2$\times$ 16.40 38.56 0.9629 34.59 0.9267 32.58 0.9048 34.30 0.9452 40.24 0.9808
EDT [49] 2$\times$ 11.48 38.63 0.9632 34.80 0.9273 32.62 0.9052 34.27 0.9456 40.37 0.9811
SemanIR-S (Ours) 2$\times$ 11.87 38.57 0.9651 34.99 0.9300 32.65 0.9078 34.86 0.9472 40.45 0.9824
SemanIR-B (Ours) 2$\times$ 19.90 38.61 0.9654 35.08 0.9304 32.69 0.9084 34.99 0.9455 40.59 0.9830
RCAN [111] 3$\times$ 15.63 34.74 0.9299 30.65 0.8482 29.32 0.8111 29.09 0.8702 34.44 0.9499
SAN [19] 3$\times$ 15.90 34.75 0.9300 30.59 0.8476 29.33 0.8112 28.93 0.8671 34.30 0.9494
HAN [68] 3$\times$ 64.35 34.75 0.9299 30.67 0.8483 29.32 0.8110 29.10 0.8705 34.48 0.9500
NLSA [64] 3$\times$ 45.58 34.85 0.9306 30.70 0.8485 29.34 0.8117 29.25 0.8726 34.57 0.9508
IPT [10] 3$\times$ 115.67 34.81 - 30.85 - 29.38 - 29.49 - - -
SwinIR [52] 3$\times$ 11.94 34.97 0.9318 30.93 0.8534 29.46 0.8145 29.75 0.8826 35.12 0.9537
CAT-A [16] 3$\times$ 16.64 35.06 0.9326 31.04 0.8538 29.52 0.8160 30.12 0.8862 35.38 0.9546
ART [102] 3$\times$ 16.58 35.07 0.9325 31.02 0.8541 29.51 0.8159 30.10 0.8871 35.39 0.9548
EDT [49] 3$\times$ 11.66 35.13 0.9328 31.09 0.8553 29.53 0.8165 30.07 0.8863 35.47 0.9550
SemanIR-S (Ours) 3$\times$ 12.05 34.99 0.9366 31.23 0.8594 29.53 0.8223 30.71 0.8950 35.52 0.9573
SemanIR-B (Ours) 3$\times$ 20.08 35.03 0.9371 31.29 0.8603 29.54 0.8227 30.87 0.9012 35.60 0.9581
RCAN [111] 4$\times$ 15.59 32.63 0.9002 28.87 0.7889 27.77 0.7436 26.82 0.8087 31.22 0.9173
SAN [19] 4$\times$ 15.86 32.64 0.9003 28.92 0.7888 27.78 0.7436 26.79 0.8068 31.18 0.9169
HAN [68] 4$\times$ 64.20 32.64 0.9002 28.90 0.7890 27.80 0.7442 26.85 0.8094 31.42 0.9177
NLSA [64] 4$\times$ 44.99 32.59 0.9000 28.87 0.7891 27.78 0.7444 26.96 0.8109 31.27 0.9184
IPT [10] 4$\times$ 115.63 32.64 - 29.01 - 27.82 - 27.26 - - -
SwinIR [52] 4$\times$ 11.90 32.92 0.9044 29.09 0.7950 27.92 0.7489 27.45 0.8254 32.03 0.9260
CAT-A [16] 4$\times$ 16.60 33.08 0.9052 29.18 0.7960 27.99 0.7510 27.89 0.8339 32.39 0.9285
ART [102] 4$\times$ 16.55 33.04 0.9051 29.16 0.7958 27.97 0.751 27.77 0.8321 32.31 0.9283
EDT [49] 4$\times$ 11.63 33.06 0.9055 29.23 0.7971 27.99 0.7510 27.75 0.8317 32.39 0.9283
SemanIR-S (Ours) 4$\times$ 12.02 33.02 0.9082 29.29 0.8026 27.96 0.7582 28.34 0.8467 32.48 0.9322
SemanIR-B (Ours) 4$\times$ 20.04 33.08 0.9090 29.34 0.8037 27.98 0.7599 28.51 0.8467 32.56 0.9335
The results of the image SR in terms of hyper-parameters $k$ under different training settings during inference for color images with different scale factors (i.e., 2$\times$, 3$\times$, and 4$\times$) are also provided in Fig. 12. All experiments are conducted in various datasets (i.e., BSD100, Manga109, Set5, Set14, and Urban100). It shows that for datasets like BSD100, Set14, and Urban100, a similar trend can be also observed in Fig. 12 compared to JPEG CAR and image denoising tasks, i.e., the random top-k setting performs more stable regardless the change of the K during the inference. However, for Manga109 and Set5 dataset. The best PSNR is obtained by the fixed top-k setting (in (d) - (i) in Fig. 12).
In general, based on all the experimental results mentioned above, we conclude that (1) the random top-k setting performs better than the fixed $k$ setting, and usually outperforms the latter by a large margin when $k$ is fixed to small values (i.e., 64, 128, 192, or 256.). (2) For the fixed top-k setting, if $k$ is set to big enough (i.e., 512) during inference, the fixed top-k setting can also achieve comparative performance or even better performance compared to the random top-k setting for several experiments (e.g.color JPEG CAR in Fig. 7 (g) and Fig. 7 (k)). However, it is not always possible that the large fixed $k$ setting can be generalized to limited computation resources, and the model trained with large fixed $k$ usually needs the same $k$ for inference to maintain the performance, which leads to heavy computation resources needed even for inference.
To this end, we propose to decouple the way to use $k$ between training and inference. i.e., we can use the random sample $k$ during training while an optional fixed $k$ during inference without degenerating the overall performance. It makes it possible to deploy models that heavily rely on large GPU memory during training but to limited GPU resources while maintaining reliable performance during inference. This is also consistent with the way we implement the proposed attention block (i.e., we adopt a Torch-Mask version that requires affordable large GPU memories during training compared to Torch-Gather while adopting the Triton version during inference) of SemanIR.
In addition, setting a predetermined $k$ value for each patch/pixel enhances computational efficiency. A fixed $k$ value facilitates parallel computation, particularly in attention operations. Conversely, making $k$ values learnable for each patch or pixel would significantly increase the complexity of the attention operation. Nonetheless, exploring the potential of learnable $k$ values for each patch or pixel represents an intriguing avenue for further investigation.
Figure 13: Visual comparison for restoring images in AWC. Best viewed by zooming.
Appendix E Experimental Results
E.1 Full Quantitative Results
Due to the limited space in the main manuscript, we only report a part of the experimental result. We show the quantitative experimental of grayscale image JPEG CAR in Tab. 11. The full results of image SR in Tab. 12, individually. These quantitative comparisons further validate the effectiveness of the proposed SemanIR.
E.2 More visual Results
To further support the effectiveness of the proposed SemanIR intuitively. We provide more visual comparison in terms of image deblurring, JPEG CAR, image denoising, and image SR below.
Image Deblurring The visual results for single image motion deblurring are shown in Fig. 14. As shown in this figure, the proposed method can effectively remove the motion blur in the input images and restore more details such as the facial contour, and the characters compared to MPRNet or Restormer.
Figure 14: Visual comparison with single image motion deblurring on GoPro dataset. Best viewed by zooming.
JPEG Compression Artifact Removal For JPEG compression artifact removal, the visual results for grayscale and color images on the Urban100 dataset are shown in Fig. 15 and Fig. 16, respectively. The proposed method achieves state-of-the-art performance in removing the blocking artifacts in the input images.
Figure 15: Visual comparison of grayscale JPEG CAR on Urban100 dataset. Best viewed by zooming.
Figure 16: Visual comparison of color JPEG CAR on Urban100 dataset. Best viewed by zooming.
Image Denoising The qualitative results for image denoising on the BSD68 and the Urban100 dataset are shown in Fig. 17 (grayscale image) and Fig. 18 (color image). It is clear that for both the grayscale and color inputs, the proposed SemanIR can remove the noise in the noisy input images and recover more realistic textural details in the restored images.
Figure 17: Visual comparison with image denoising on BSD68 dataset. Best viewed by zooming.
Figure 18: Visual comparison with image denoising on Urban100 dataset. Best viewed by zooming.
IR Adverse Weather Conditions. The qualitative results for IR in AWC on the Test1 [48, 47] dataset are shown in Fig. 13. It shows a challenging case but our method can restore better structural content and clearer details.
Image SR. The comparison of visual results of different image SR methods is shown in Fig. 19 and Fig. 20. Fig. 19 shows the results on the Urban100 dataset, and Fig. 20 shows the results on the Manga109 dataset. The proposed SemanIR can restore more missing details in the LR images compared to other state-of-the-art methods like SwinIR, ART, CAT, and EDT.
Figure 19: Visual comparison (4$\times$) with image SR on Urban100 dataset. Best viewed by zooming.
Figure 20: Visual comparison (4$\times$) with image SR on Manga109 dataset. Best viewed by zooming.